OpenStudy (anonymous):

5x + 2y = 8 x + y = 4 If you want to solve the system of equations by addition, which of the following could you do? Multiply the second equation by 5 and add. Multiply the second equation by 2 and add. Multiply the second equation by -2 and add.

OpenStudy (anonymous):

Multiply the second equation by -2 and add.

OpenStudy (wolf1728):

Yes multiply x + y = 4 by -2 then add So we have -2x -2y = -8 which we add to 5x + 2y = 8

OpenStudy (wolf1728):

Which produces x =0 Now, can you solve for y?